a) Escape velocity from the Moon is approximately 5500 miles per hour = 2.45 x 103m/s. At certain spots on the surface of the Moon, the temperature of the surface material reaches 800K. Suppose a molecule of hydrogen (m= 3.33x 0-27kg) were to hit the surface of the Moon and reach a speed equal to the root-mean-square (rms) speed for this temperature. `
1
Expert's answer
2020-11-06T10:13:52-0500
"v=\sqrt{\frac{3RT}{M}}=\sqrt{\frac{3(8.31)(800)}{10^{-3}}}=4466\frac{m}{s}>2450\frac{m}{s}"